OpenStudy (cresciez):

Medal and fan to anyone who knows how to do this! You buy equal numbers of containers of tomatoes, cucumbers, and carrots. Both carrots and cucumbers are $3 less per container than tomatoes. The product of the price for each equals the total cost. What is the price for each container? Write a polynomial function to describe the cost of your purchase. Graph the function. (Let: t = cost for each tomato container)

OpenStudy (cresciez):

Answers:

OpenStudy (cresciez):

Just a side note I actually have no clue on how to do this at all. Ahah oops

OpenStudy (tkhunny):

Well, to be honest, the problem makes no sense. The Total Price is the SUM of the individual prices, not the PRODUCT. That's why I'll just suggest to you how it might be answered, even though the problem makes no sense at all. t = cost for tomato container. Based on this information: "Both carrots and cucumbers are $3 less per container than tomatoes." We have" t - 3 = Cost for carrots t - 3 = Cost for cukes From the silly "Product" idea, we have Price(t) = t(t-3)(t-3) = t(t-3)^2 This is a cubic, not a quadratic. This realization discards Choice D and probably discards Choice A. It should also be defined only for Positive numbers or zero for "t". This also discards Choice D. It should not EVER give a Price that is negative. This discards Choice B. It should pass through the Origin. That's what the "t" does in the formula. This discards the now-beleaguered Choice D AGAIN. It should bounce off the horizontal axis at t = 3. That's what (t-3)^2 does in the formula. This final fact eliminates all but Choice C. And there we have it.

OpenStudy (cresciez):

Thank you so much for your help!